# A phase-space view of vibrational energies without the Born-Oppenheimer framework

### v1

@@ I.2 Review of Electronic Phase-Space Hamiltonians
$\Big{[}-i\hbar\sum_{B}\left(\bm{R}_{B}\times\frac{\partial{}}{\partial{\bm{R}_{B}}}\right)_{\gamma}+\hat{l}_{\gamma}+\hat{s}_{\gamma},\hat{\Gamma}_{A\delta}\Big{]}$
$=i\hbar\sum_{\alpha}\epsilon_{\alpha\gamma\delta}\hat{\Gamma}_{A\alpha},$ (22)
and thus we may conclude that the $\hat{\bm{\Gamma}}^{\prime}$ and $\hat{\bm{\Gamma}}^{\prime\prime}$ operators are akin to the local translational and rotational components of the derivative couplings[24, 25, 26, 27, 23], so that the phase-space Hamiltonian in Eq. 14 is an approximate (but tractable) version of Eq. 17.
I.3 Our Goal In this Article
@@
This concludes our heuristic review of phase-space electronic Hamiltonian theory. We have sought to justify the phase-space approach by appealing to the readers’ understanding of the importance of a change of coordinates when simulating a simple hydrogen atom, but have also demonstrated that the approach is reasonable by mapping the approximate $\hat{\bm{\Gamma}}$ factors to the derivative coupling. By this point in time, there is already ample evidence in the literature that this phase-space approach can recover electronic momenta[22], electronic current densities[23] and even vibrational circular dichroism (VCD) signals[28]. In all cases, we have found that a phase-space electronic Hamiltonian is simply more accurate, and offers more and richer set of information than a BO Hamiltonian.
@@
That being said, however, one could argue that we still do not have a smoking gun that clearly establishes the supremacy of the phase-space approach. Furthermore, one might also note that the phase-space approach developed in Sec. I.2 does require some hand-waving, and for complete confidence, one would like to understand how to derive and work with such a Hamiltonian in the context of an exact calculation. These concerns motivate the present article. In what follows, we will first establish how a phase-space electronic Hamiltonian can be rigorously defined through a Wigner transformation, and then, using the inverse Weyl transformation, we will show how to extract quantized vibrational energies from a phase-space approach. In particular, we will show that vibrational energies extracted by diagonalizing the nuclear Hamiltonian from a phase-space electronic Hamiltonian universally outperforms (i.e. gives more accurate vibrational energies) than diagonalizing the corresponding nuclear Hamiltonian from a Born-Oppenheimer electronic Hamiltonian. This finding would appear to be the smoking gun that makes clear that, in the future, quantum chemistry packages will need to be prepared to avoid the Born-Oppenheimer Hamiltonian.
II Born-Oppenheimer versus phase-space theory of vibrational energies
@@
Before introducing phase-space theory, it will be helpful to review briefly the easiest way to derive BO theory from a formal point of view.
II.1 Standard BO theory
@@
The BO assumption is then that one can neglect the $\hat{\bm{P}}\cdot\hat{\bm{d}}$ and $\hat{\bm{d}}\cdot\hat{\bm{d}}$ terms in Eq. 17, and diagonalize:
@@
Note that, because $\hat{V}_{\rm ad}={\rm diag},(\epsilon_{1}(\hat{\bm{R}}),\cdots,\epsilon_{n}(\hat{\bm{R}}))$ is diagonal, this amounts to many different diagonalizations of $\frac{\hat{\bm{P}}^{2}}{2M}+\epsilon_{n}({\bm{R}})$. Note also that, while $\hat{H}_{\rm ad}$ is a unitary transform of $\hat{H}$, $\hat{H}_{\rm BO}$ is clearly not such a transform.
Everything above is well known and standard. Also well known, but slightly less standard, is that the BO approximation can also be derived through partial Wigner transforms[29, 30, 31]. As a reminder, the partial Wigner transform ($\mathcal{W}$) for any operator $\hat{O}$ is given by:
@@
where $\hat{O}_{W}({\bm{R}},{\bm{P}})$ is an matrix of electronic operators that depends parametrically on ${\bm{R}}$ and ${\bm{P}}$. This transformation is invertible by a Weyl transform, and one can show that:
@@
$\bra{\bm{R}}\hat{O}\ket{\bm{R}^{\prime}}$ $=$ $\int\frac{d\bm{P}}{2\pi\hbar}e^{\frac{i}{\hbar}\bm{P}\cdot(\bm{R}-\bm{R}^{\prime})}\hat{O}_{W}\left(\frac{\bm{R}+\bm{R}^{\prime}}{2},P\right).$ (28)
Now, to formally derive the BO approximation, we build and diagonalize the Wigner transform of $\hat{H}$:
@@
The BO Hamiltonian in Eq. 24 is then exactly equivalent to the Weyl transform of $\Lambda_{W}$:
@@
Note that the BO Hamiltonian is (obviously) not a unitary transform of the exact Hamiltonian. If one were to apply a Wigner transformation to the adiabatic Hamiltonian in Eq. 23, one must remember that the Wigner transform of a product of operators is not the product of Wigner transforms but rather involves a Moyal product between them. In such a case, it follows that the exact Hamiltonian can be written as:
@@
Observe that the difference between Eq. 30 and Eq. 31 is that normal product multiplication is replaced by Moyal star products.
In any event, it is quite common today to extract vibrational energies from a BO Hamiltonian. To do so, for the ground state, one simply diagonalizes the nuclear Hamiltonian for the electronic ground state $\bra{\psi_{0}}\hat{H}_{\rm BO}\ket{\psi_{0}}$ (which is equivalent to diagonalizing $\frac{\hat{\bm{P}}^{2}}{2M}+\epsilon_{0}(\hat{\bm{R}})$). Alternatively, within a harmonic approximation, one can can simply expand $\epsilon_{0}$ (or equivalently $\bra{\psi_{0}}\hat{\Lambda}_{W}\ket{\psi_{0}}$) to second order in $\bm{R}$ and apply the standard harmonic oscillator result[32].
II.2 Phase-Space Electronic Hamiltonian
According the theory presented in Sec. I.2, there are better means to approximate the full Hamiltonian than $\hat{H}_{\rm BO}$. In generating such approximations, and in light of Eq. 16, it would appear that the optimal approach is to parameterize the approximate Hamiltonian by both nuclear position and momentum. Rigorously, such a parameterization can only occur after a Wigner transformation. Thus, the phase-space approach is to construct an electronic Hamiltonian of the form:
@@ II.2 Phase-Space Electronic Hamiltonian
The final phase-space electronic Hamiltonian is then constructed (as for BO) by a Weyl transformation:
@@
Eq. 34 should be compared with Eq. 30. Note that, just like $\hat{H}_{\rm BO}$, $\hat{H}_{\rm PS}$ is not a unitary transform of $\hat{H}$. Also, just as for $H_{\rm BO}$, vibrational energies can be formally extracted by diagonalizing the nuclear Hamiltonian $\bra{\psi_{0}^{\rm PS}}\hat{H}_{\rm PS}\ket{\psi_{0}^{\rm PS}}$, but in this case, constructing $\hat{H}_{\rm PS}$ requires a Weyl transform in Eq. 34. Put differently, while working with the phase-space electronic Hamiltonian is obvious for semiclassical simulations, for quantum nuclear (i.e. vibrational) problems, where ${\bm{P}}$ is an operator (not a paramter), the phase-space electronic Hamiltonian has no simple analogue to Eq. 24 (but rather requires a Weyl transformation).
@@
III Demonstration through a model problem
To demonstrate the power of a phase-space approach for calculating vibrational spectra, consider a classic model Hamiltonian that describes a one-dimensional asymmetric hydrogen bond system $\rm O-H\cdots O$ from Refs.33, 14. In this coarse-grained model, the light particle represents a hydrogen atom with mass $m$, and the heavy particles are oxygen atoms with mass $M$ – though we will also refer to the particle as an electron and the heavy particle as a nucleus, given the importance of BO theory to electronic structure calculations. The potential from Ref.33, 14 is plotted in Fig. 1(a) as a function of the heavy and light degrees of freedom (DoF) $R$ and $r$, where $R$ is the distance between the two oxygen atoms and the light DoF $r$ indicates the distance between the H atom to the center of two O atoms. The intuition behind this model Hamiltonian is that the hydrogen will stay at the center when the two oxygen atoms are close, and it will bonded to one oxygen when the oxygen atoms are far away due to the hydrogen bonding effects. All of the relevant details of the Hamiltonian can be found in SI.

### v2

@@ 2.2 Phase-Space Electronic Hamiltonians and a Simple Form of $\hat{\bm{\Gamma}}_{A}$
$\Big{[}-i\hbar\sum_{B}\left(\bm{R}_{B}\times\frac{\partial{}}{\partial{\bm{R}_{B}}}\right)_{\gamma}+\hat{l}_{\gamma}+\hat{s}_{\gamma},\hat{\Gamma}_{A\delta}\Big{]}$ $=$ $i\hbar\sum_{\alpha}\epsilon_{\alpha\gamma\delta}\hat{\Gamma}_{A\alpha},$ (38)
Moreover, one can show that dynamics along an eigenstate of a Hamiltonian of the form in Eq. 34 (where $\hat{\bm{\Gamma}}$ satisfies Eqs. 35-38) will conserve linear and angular momentum22, 23, 24.
Alas, $\hat{\bm{\Gamma}^{\prime}}$ in Eq. 28 satisfies only Eqs. 35 and 36, but not Eqs. 37 and 38. The implication of this fact is that, if one diagonalizes Eq. 27, then classical dynamics propagated along the resulting energy surfaces will conserve linear momentum24 but not angular momentum. Nevertheless, we have argued that one can use $\hat{\bm{\Gamma}}^{\prime}$ in Eq. 28 as a starting point, and build a proper $\hat{\bm{\Gamma}}$ as the sum of two contributions:
$\hat{\bm{\Gamma}}=\hat{\bm{\Gamma}}^{\prime}+\hat{\bm{\Gamma}}^{\prime\prime}$ (39)
with final form of the phase-space Hamiltonian being 22:
All that remains is to build the $\hat{\bm{\Gamma}}^{\prime}$ and $\hat{\bm{\Gamma}}^{\prime\prime}$ operators, which can be thought of as the local translational and rotational components of the derivative couplings25, 26, 27, 28, 24. Let us now describe how to construct these objects explicitly.
2.2.1 Form of $\hat{\bm{\Gamma}}_{A}^{\prime}$
For the construction of $\hat{\bm{\Gamma}^{\prime}}$ in Eq. 28, it remains only to construct the partition of unity function $\hat{\theta}$. We have chosen to define $\hat{\theta}_{A}$ as combinations of Gaussian functions localized on different atoms:
$\hat{\theta}_{A}(\hat{\bm{r}})=\frac{M_{A}e^{-|\hat{\bm{r}}-\bm{R}_{A}|^{2}/\sigma_{A}^{2}}}{\sum_{B}M_{B}e^{-|\hat{\bm{r}}-\bm{R}_{B}|^{2}/\sigma_{B}^{2}}}.$ (41)
In Eq. 41, we utilize Gaussian functions with width $\sigma_{A}$, centered around $\bm{R}_{A}$ and weighted by $M_{A}$. The physical meaning of this spatial partition function $\hat{\theta}_{A}$ is then clear; $\hat{\theta}_{A}$ measures the extent to which the electrons are influenced by nucleus $A$, i.e., electrons are more strongly affected by nucleus $A$’s motion when they are closer to $A$ or $A$ is heavier. Note that the parameter $\sigma_{A}$ determines the locality of this measure. In practical ab initio electronic structure calculations, as discussed in Ref. 24, an optimal choice for $\sigma_{A}$ can be $\sigma_{A}=0.2r_{\rm vdW}^{A}$, where $r_{\rm vdW}^{A}$ is the van der Waals radius of atom $A$.
2.2.2 Form of $\hat{\bm{\Gamma}}_{A}^{\prime\prime}$
For the rotational component $\hat{\bm{\Gamma}}_{A}^{\prime\prime}$, we define the $\hat{\bm{\Gamma}}_{A}^{\prime\prime}$ to be the basis-free electronic rotation factor:
$\hat{\bm{\Gamma}}_{A}^{\prime\prime}$ $=$ $\sum_{B}\zeta_{AB}\left(\bm{R}_{A}-\bm{R}^{0}_{B}\right)\times\left(\bm{K}_{B}^{-1}\hat{\bm{J}}_{B}\right),$ (42)
where
$\hat{\bm{J}}_{B}$ $=$ $\frac{1}{2i\hbar}\left((\hat{\bm{r}}-\bm{R}_{B})\times\hat{\theta}_{B}(\hat{\bm{r}})\hat{\bm{p}}+(\hat{\bm{r}}-\bm{R}_{B})\times\hat{\bm{p}}\hat{\theta}_{B}(\hat{\bm{r}})+2\hat{\bm{s}}\hat{\theta}_{B}\right),$ (43)
$\bm{R}_{B}^{0}$ $=$ $\frac{\sum_{A}\zeta_{AB}\bm{R}_{A}}{\sum_{A}\zeta_{AB}},$ (44)
$\bm{K}_{B}$ $=$ $\sum_{A}\zeta_{AB}\left(\bm{R}_{A}\bm{R}_{A}^{\top}-\bm{R}_{B}^{0}\bm{R}_{B}^{0\top}-(\bm{R}_{A}^{\top}\bm{R}_{A}-\bm{R}_{B}^{0\top}\bm{R}_{B}^{0})\mathcal{I}_{3}\right),$ (45)
and the symmetric localization function $\zeta_{AB}$
$\zeta_{AB}$ $=e^{-\frac{|\bm{R}_{A}-\bm{R}_{B}|^{2}}{2(\sigma_{A}+\sigma_{B})^{2}}}.$ (46)
It is straightforward to verify that the $\hat{\bm{\Gamma}}$ term defined above satisfies the following four constraints in Eqs. 35-3824. Altogether, Eqs. 40-46 constitute a complete, self-contained, practical phase-space Hamiltonian $\hat{H}_{\rm PS}$.
2.3 Phase-Space Electronic Hamiltonian and Its Connection to Berry’s Superadiabatic Formalism
At this point, we have sought to justify the use of an electronic Hamiltonian by noting that the approach includes some non-Born Oppenheimer effects by approximating the derivative coupling ($\hat{\bm{d}}$) with an operator ($\hat{\bm{\Gamma}}$). That being said, it is important to note that a similar conclusion can also be drawn by considering Berry’s superadiabatic formalism–which is not very well known in chemistry and which we now briefly review.
Consider a Hamiltonian $\hat{H}(\lambda(t))$ with a slow varying time-dependent parameter $\lambda$. Let us diagonalize the Hamiltonian to obtain the zeroth-order adiabatic basis {|$\{\ket{\psi^{0}_{j}}\}$,
$\hat{H}\ket{\psi^{0}_{j}}=E^{0}_{j}\ket{\psi^{0}_{j}}.$ (47)
and expand the system wavefunction in this adiabatic basis,
|$\ket{\Psi}=\sum_{j}c_{j}^{0}\ket{\psi^{0}_{j}}.$ (48)
In this basis, the time-dependent Schrodinger equation can be expressed as
$\dot{c}_{j}^{0}=-\frac{i}{\hbar}E_{j}^{0}c_{j}^{0}+\sum_{k}\dot{\lambda}\bra{\psi_{j}^{0}}\nabla_{\lambda}\ket{\psi_{k}^{0}}c_{k}^{0}.$ (49)
Thus, an effective first-order Hamiltonian can be identified from Eq. 49 as:
$H^{1}_{jk}=E_{j}^{0}-i\hbar\dot{\lambda}\bra{\psi_{j}^{0}}\nabla_{\lambda}\ket{\psi_{k}^{0}},$ (50)
As noted by Ref. 29, one can diagonalize this Hamiltonian to generate a first-order super-adiabatic basis
$\hat{H}^{1}\ket{\psi_{j}^{1}}=E_{j}^{1}\ket{\psi_{j}^{1}}.$ (51)
Moreover, as the first-order superadiabatic basis remains time-dependent, this diagonalization process can be applied iteratively to construct higher and higher-order superadiabatic bases.
Now, the hope is that by performing such iterative diagonalizations, one can arrive at a superadiabatic basis, whereby non-adiabatic effects can be effectively suppressed (even if it will not completely vanish), making the propagation of adiabatic dynamics along superadiabatic surface more accurate. Moreover, in the context of semiclassical molecular dynamics, in a landmark paper30, Shenvi showed that, to first order only, this approach can be used to generate phase-space electronic Hamiltonians of the form:
$H_{{\rm PS},mn}^{\rm Shenvi}(\bm{R},\bm{P})=\sum_{k}\frac{({\bm{P}}_{A}\delta_{mk}-i\hbar{\bm{d}}^{A}_{mk})({\bm{P}}_{A}\delta_{kn}-i\hbar{\bm{d}}^{A}_{kn})}{2M_{A}}+E_{n}^{\rm BO}({\bm{R}})\delta_{mn}$ (52)
Shenvi posited that surface hopping dynamics along these surfaces can be used for so-called phase-space surface hopping (PSSH) dynamics, and Ref.30 demonstrates that for some particular cases, PSSH can vastly outperform more standard fewest switches surface hopping dynamics (FSSH)31.
At this point, however, it is worth noting that there are several problems with Eq. 52. In particular, this Hamiltonian becomes problematic at the vicinity of conical intersections, where $\bm{d}_{jk}$ diverges. This Hamiltonian is also computationally expensive, as calculating the forces on the superadiabatic surfaces would appear to require evaluating second-order derivatives (the derivatives of derivative couplings). Finally, treating degenerate adiabatic states (with potentially arbitrary gauge choices) becomes unreliable; in such a case, the derivative couplings are not even well-defined. For all of the reasons above, very little work has been achieved using an electronic phase-space Hamiltonian of the form in Eq. 52 beyond model problems32. Nevertheless, as we have argued above, if one can approximate $\hat{\bm{d}}$ by constructing a meaningful (and tractable) $\hat{\bm{\Gamma}}$ operator, we believe there is a strong future for electronic phase-space Hamiltonian of the form in Eq. 40.
2.4 Our Goal In this Article
@@
This concludes our heuristic review of phase-space electronic Hamiltonian theory. We have sought to justify a phase-space approach by appealing to the readers’ understanding of the importance of a change of coordinates when simulating a simple hydrogen atom, but have also demonstrated that the approach is reasonable by mapping the approximate $\hat{\bm{\Gamma}}$ factors to the derivative coupling. By this point in time, there is already ample evidence in the literature that this phase-space approach can recover electronic momenta23, electronic current densities24 and even vibrational circular dichroism (VCD) signals33. In all cases, we have found that a phase-space electronic Hamiltonian is simply more accurate, and offers more and richer set of information than a BO Hamiltonian.
@@
That being said, however, one could argue that we still do not have a smoking gun that clearly establishes the supremacy of the phase-space approach. Furthermore, one might also note that the phase-space approach developed in Sec. 2 does require some hand-waving, and for complete confidence, one would like to understand how to derive and work with such a Hamiltonian in the context of an exact calculation. These concerns motivate the present article. In what follows, we will first establish how a phase-space electronic Hamiltonian can be rigorously defined through a Wigner transformation, and then, using the inverse Weyl transformation, we will show how to extract quantized vibrational energies from a phase-space approach. In particular, we will show that vibrational energies extracted by diagonalizing the nuclear Hamiltonian from a phase-space electronic Hamiltonian universally outperforms (i.e. gives more accurate vibrational energies) than diagonalizing the corresponding nuclear Hamiltonian from a Born-Oppenheimer electronic Hamiltonian. This finding would appear to be the smoking gun that makes clear that, in the future, quantum chemistry packages will need to be prepared to avoid the Born-Oppenheimer Hamiltonian.
3 Born-Oppenheimer vibrational energies
@@
Before demonstrating how to extract vibrational energies from phase-space theory, it will be helpful to review briefly the easiest way to derive BO theory and BO vibrational energies from a formal point of view for a system of many electrons and nuclei (and thus expand on the simple approach for the hydrogen atom in Sec. 1 above).
@@
The BO approximation is then that one can neglect the $\hat{\bm{P}}\cdot\hat{\bm{d}}$ and $\hat{\bm{d}}\cdot\hat{\bm{d}}$ terms in Eq. 53, and diagonalize:
@@
Note that, because $\hat{V}_{\rm ad}={\rm diag}(E_{0}(\hat{\bm{R}}),\cdots,E_{n}(\hat{\bm{R}}))$, the BO assumption amounts to many different diagonalizations of $\frac{\hat{\bm{P}}^{2}}{2M}+E_{n}({\bm{R}})$. Note also that, while $\hat{H}_{\rm ad}$ is a unitary transform of $\hat{H}$, $\hat{H}_{\rm BO}$ is clearly not such a transform.
Everything above is well known and standard. Also well known, but slightly less standard, is that the BO approximation can also be derived through partial Wigner transforms34, 35, 36. As a reminder, the partial Wigner transform ($\mathcal{W}$) for any operator $\hat{O}$ is given by:
@@
where $\hat{O}_{W}({\bm{R}},{\bm{P}})$ is a matrix of electronic operators that depends parametrically on ${\bm{R}}$ and ${\bm{P}}$. This transformation is invertible by a Weyl transform, and one can show that:
@@
$\bra{\bm{R}}\hat{O}\ket{\bm{R}^{\prime}}$ $=$ $\int\frac{d\bm{P}}{2\pi\hbar}e^{\frac{i}{\hbar}\bm{P}\cdot(\bm{R}-\bm{R}^{\prime})}\hat{O}_{W}\left(\frac{\bm{R}+\bm{R}^{\prime}}{2},\bm{P}\right).$ (58)
Now, to formally derive the BO approximation, we build and diagonalize the partial Wigner transform of $\hat{H}$:
@@
Note that, on the right hand side of Eq. 59, $\hat{U}_{W}$ is an electronic unitary operator, parametrized by nuclear position ${\bm{R}}$ and we perform electronic matrix multiplication. The BO Hamiltonian in Eq. 54 is then exactly equivalent to the Weyl transform of $\Lambda_{W}$:
@@
Note that the BO Hamiltonian is (obviously) not a unitary transform of the exact Hamiltonian. If one were to apply a partial Wigner transformation to the adiabatic Hamiltonian in Eq. 53, one must remember that the Wigner transform of a product of operators is not the product of Wigner transforms but rather involves a Moyal product between them:
$\left(\hat{A}\hat{B}\right)_{W}=A_{W}*B_{W}.$ (61)
Here, the Moyal product between two phase-space functions $A_{W}(\bm{R},\bm{P})$ and $B_{W}(\bm{R},\bm{P})$ is defined as
$A_{W}(\bm{R},\bm{P})*B_{W}(\bm{R},\bm{P})=A_{W}(\bm{R},\bm{P})e^{-\frac{i\hbar}{2}\hat{\Lambda}}B_{W}(\bm{R},\bm{P})$ (62)
where the Possion bracket operator $\hat{\Lambda}=\reflectbox{$\vec{\reflectbox{$\bm{\nabla}$}}$}_{\bm{P}}\cdot\vec{\bm{\nabla}}_{\bm{R}}-\reflectbox{$\vec{\reflectbox{$\bm{\nabla}$}}$}_{\bm{R}}\cdot\vec{\bm{\nabla}}_{\bm{P}}$. In such a case, it follows that the exact Hamiltonian can be written as:
@@
Observe that the difference between Eq. 60 and Eq. 63 is that normal product multiplication is replaced by Moyal products.
In any event, it is quite common today to extract vibrational energies from a BO Hamiltonian. To do so, for the ground state, one simply diagonalizes the nuclear Hamiltonian for the electronic ground state $\bra{\psi_{0}}\hat{H}_{\rm BO}\ket{\psi_{0}}$ (which is equivalent to diagonalizing $\frac{\hat{\bm{P}}^{2}}{2M}+E_{0}(\hat{\bm{R}})$). Alternatively, within a harmonic approximation, one can can simply expand $E_{0}$ (or equivalently $\bra{\psi_{0}}\hat{\Lambda}_{W}\ket{\psi_{0}}$) to second order in $\bm{R}$ and apply the standard harmonic oscillator result37.
4 Vibrational Energies from a Phase-Space Electronic Hamiltonian
According to the theory presented in Sec. 2, there are better means to approximate the full Hamiltonian than $\hat{H}_{\rm BO}$. In generating such approximations, and in light of Eq. 40, it would appear that the optimal approach is to parameterize the approximate Hamiltonian by both nuclear position and momentum. Rigorously, such a parameterization can only occur after a Wigner transformation. Thus, the phase-space approach is to construct an electronic Hamiltonian of the form:
@@ 4 Vibrational Energies from a Phase-Space Electronic Hamiltonian
As above for the BO case, on the right hand side of Eq. 65, $\hat{L}_{W}$ is an electronic unitary operator – but now parametrized by both nuclear position ${\bm{R}}$ and nuclear momentum ${\bm{P}}$ – and we perform electronic matrix multiplication in Eq. 65. The final phase-space electronic Hamiltonian is then constructed (as for BO) by a Weyl transformation:
@@
Eq. 66 should be compared with Eq. 60. Note that, just like $\hat{H}_{\rm BO}$, $\hat{H}_{\rm PS}$ is not a unitary transform of $\hat{H}$. Also, just as for $H_{\rm BO}$, vibrational energies can be formally extracted by diagonalizing the nuclear Hamiltonian $\bra{\psi_{0}^{\rm PS}}\hat{H}_{\rm PS}\ket{\psi_{0}^{\rm PS}}$; however, unlike the BO case, constructing $\hat{H}_{\rm PS}$ requires a Weyl transform in Eq. 66. Put differently, while working with the phase-space electronic Hamiltonian is obvious for semiclassical simulations, for quantum nuclear (i.e. vibrational) problems, where ${\bm{P}}$ is an operator (not a paramter), the phase-space electronic Hamiltonian has no simple analogue to Eq. 54 (but rather requires a Weyl transformation).
@@
Before concluding this section, one item must be emphasized. For the reader who is scared of Wigner transforms, note that a harmonic phase-space vibrational energy (which can be calculated by a quadratic expansion and entirely without a Wigner transformation) always outperforms a harmonic BO calculation for vibrational energies. As such, there is not only a conceptual but also a very practical reason to employ a phase-space Hamiltonian in the immediate future.
5 Demonstration through a model problem
To demonstrate the power of a phase-space approach for calculating vibrational spectra, consider a classic model Hamiltonian that describes a one-dimensional asymmetric hydrogen bond system $\rm O-H\cdots O$ from Refs.38, 13. In this coarse-grained model, the light particle represents a hydrogen atom with mass $m$, and the heavy particles are oxygen atoms with mass $M$ – though we will also refer to the particle as an electron and the heavy particle as a nucleus, given the importance of BO theory to electronic structure calculations.
5.1 Model Hamiltonian
The Hamiltonian of the three-particle one-dimension hydrogen bond model can be written as:
$\hat{H}=\frac{\hat{P}_{1}^{2}}{2M}+\frac{\hat{P}_{2}^{2}}{2M}+\frac{\hat{p}^{2}}{2m}+\hat{V}(\hat{R}_{1},\hat{R}_{2},\hat{r}_{\rm e}),$ (67)
where we assume the mass of two heavy particles are the same $M_{1}=M_{2}=M$. We further reduce the complexity of this Hamiltonian by going to the molecular (nuclei + electron) center of mass frame. Namely, we define the molecular center of mass (MCM)
$\hat{R}_{\rm MCM}=\frac{M\hat{R}_{1}+M\hat{R}_{2}+m\hat{r}_{\rm e}}{2M+m},$ (68)
We transform the coordinates from $(\hat{R}_{1},\hat{R}_{2},\hat{r}_{\rm e})$ to MCM coordinates $(\hat{R}_{\rm MCM},\hat{R},\hat{r})$, where $\hat{R}$ is the distance between the two oxygen atoms and the light degrees of freedom $\hat{r}$ indicates the distance between the H atom to the nuclear center of mass.
$\hat{R}=\hat{R}_{1}-\hat{R}_{2},\quad\hat{r}=\hat{r}_{\rm e}-\frac{\hat{R}_{1}+\hat{R}_{2}}{2}.$ (69)
The Hamiltonian therefore becomes effectively a two dimensional model if we get rid of the molecular center of mass motion:
$\hat{H}=\frac{\hat{P}_{\rm MCM}^{2}}{2(2M+m)}+\frac{\hat{P}^{2}}{2\mu}+\frac{{\hat{p}}^{2}}{2m}+\frac{{\hat{p}}^{2}}{2(M+M)}+\hat{V}(\hat{R},\hat{r}),$ (70)
where $\mu=\frac{M}{2}$ is the nuclear reduced mass and the second to last term is the mass polarization term, and the momentum operators in the new coordinate system are defined as:
$\hat{P}_{1}=\frac{M\hat{P}_{\rm MCM}}{2M+m}+\hat{P}-\frac{\hat{p}}{2},\quad\hat{P}_{2}=\frac{M\hat{P}_{\rm MCM}}{2M+m}-\hat{P}-\frac{\hat{p}}{2},\quad\hat{p}_{\rm e}=\hat{p}+\frac{m\hat{P}_{\rm MCM}}{2M+m}.$ (71)
In the MCM coordinates, the potential energy term does not depends on the $R_{\rm MCM}$ and can be expressed as:
$\hat{V}(\hat{R},\hat{r})$ $=$ $D\left(e^{-2a\left(\frac{\hat{R}}{2}+\hat{r}-d\right)}-2e^{-a\left(\frac{\hat{R}}{2}+\hat{r}-d\right)}+1\right)$ (72)
$+$ $Dc^{2}\left(e^{-\frac{2a}{c}\left(\frac{\hat{R}}{2}-\hat{r}-d\right)}-2e^{-\frac{a}{c}\left(\frac{\hat{R}}{2}-\hat{r}-d\right)}\right)+Ae^{-B\hat{R}}-\frac{C}{\hat{R}^{6}}.$
where the parameters used in the paper are given in Table. 1.
$D$ $d$ $a$ $c$ $A$ $B$ $C$
$60~{}\rm kcal/mol$ $0.95~{}\textup{\AA}$ $2.52~{}\textup{\AA}^{-1}$ $0.707$ $2.32\times 10^{5}~{}\rm kcal/mol$ $3.15~{}\textup{\AA}^{-1}$ $2.31\times 10^{4}~{}\rm kcal/mol/\textup{\AA}^{6}$
Table 1: Model parameters from Ref. 38.
The potential from Refs.38, 13 is plotted in Fig. 1(a) as a function of the heavy and light degrees of freedom $R$ and $r$. The intuition behind this model Hamiltonian is that the hydrogen will stay at the center when the two oxygen atoms are close, and it will bonded to one oxygen when the oxygen atoms are far away due to the hydrogen bonding effects.
5.2 Methods